OpenStudy (anonymous):

The "alum" used in cooking is potassium aluminum sulfate hydrate, KA1(SO4)2 x H2O. To find the value of x, you can heat a sample of the compound to drive off all of the water and leave only KA1(SO4)2. Assume you heat 4.74 g of the hydrated compound and that the sample loses 2.16 g of water. What is the value of x?

OpenStudy (jhannybean):

@aaronq

OpenStudy (aaronq):

mass of \(\sf KAl(SO_4)_2*xH_2O\) = 4.74 g Find moles of the anhydrate and of water, separately, using: \(\sf moles=\dfrac{mass}{Molar~mass}\) \(\sf moles~of ~ KAl(SO_4)_2=\dfrac{4.74g-2.16g}{258.21~g/mol }=0.01 ~moles\) \(\sf moles~of ~H_2O=\dfrac{2.16g}{18~g/mol}=0.12~moles\) Normalize by dividing both by the lowest number \(KAl(SO_4)_2=\dfrac{0.01}{0.01}=1\) \(H_2O=\dfrac{0.12}{0.01}=12\) the number of x is 12, the formula of the hydrate is \(\sf KAl(SO_4)_2*12~H_2O\)